### v4

@@ 1.2 Complexity Analysis
achieves a convergence rate in distribution for the NSA scheme (7)-(8) of order $\varepsilon$ as $\varepsilon\downarrow 0$. The corresponding complexity satisfies
@@
for some positive constant $C<\infty$. By Corollary 1.4, the convergence rate of NSA, with bias parameter $h$ and iterations amount $n=\lceil h^{-2}\rceil$, is of order $h^{\beta}$ as $\mathcal{H}\ni h\downarrow 0$. The result follows by setting $h$ such that $h^{\beta}\leq\varepsilon$.
@@ 2 Averaged Nested Stochastic Approximation Algorithm
According to Theorem 1.2 and Corollary 1.4, the best CLT convergence rate is achieved for $\beta=1$, i.e. by setting $\gamma_{n}=\gamma_{1}n^{-1}$ for the VaR NSA scheme. This choice is only possible under the additional constraint $\lambda\gamma_{1}>1$, where $\lambda$ is an explicit yet inaccessible constant (9). In practice, $\gamma_{1}$ must be fine-tuned, adding a computational burden to the algorithm’s usage.
@@
$\overline{\xi}^{h}_{n}=\frac{1}{n}\sum_{k=1}^{n}\xi^{h}_{k}=\Big(1-\frac{1}{n}\Big)\overline{\xi}^{h}_{n-1}+\frac{1}{n}\xi^{h}_{n},\quad n\geq 1,$ (13)
where $\overline{\xi}^{h}_{0}=0$ and $\{\xi^{h}_{n},n\geq 0\}$ are obtained via the NSA scheme (7) with step sizes $\{\gamma_{n}=\gamma_{1}n^{-\beta},n\geq 1\}$, $\gamma_{1}>0$, $\beta\in\big(\frac{1}{2},1\big)$.
@@ 2.1 Convergence Rate Analysis
$h^{-1}\begin{pmatrix}\overline{\xi}^{h}_{\lceil h^{-2}\rceil}-\xi^{h}_{\star}\\ \chi^{h}_{\lceil h^{-2}\rceil}-\chi^{h}_{\star}\end{pmatrix}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{\mathcal{L}}}}\mathcal{N}(0,\overline{\Sigma})\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0,$
@@
$\overline{\Sigma}=\begin{pmatrix}\frac{\alpha(1-\alpha)}{f_{X_{0}}(\xi^{0}_{\star})^{2}}&\frac{\alpha}{1-\alpha}\frac{\mathbb{E}[(X_{0}-\xi^{0}_{\star})^{+}]}{f_{X_{0}}(\xi^{0}_{\star})}\\ \frac{\alpha}{1-\alpha}\frac{\mathbb{E}[(X_{0}-\xi^{0}_{\star})^{+}]}{f_{X_{0}}(\xi^{0}_{\star})}&\frac{\mathbb{V}\mathrm{ar}((X_{0}-\xi^{0}_{\star})^{+})}{(1-\alpha)^{2}}\end{pmatrix}$. (14)
Proof 2.3.
See Appendix C.
@@
$h^{-1}\begin{pmatrix}\overline{\xi}^{h}_{\lceil h^{-2}\rceil}-\xi^{0}_{\star}\\ \chi^{h}_{\lceil h^{-2}\rceil}-\chi^{0}_{\star}\end{pmatrix}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{\mathcal{L}}}}\mathcal{N}\left(\begin{pmatrix}-\frac{v(\xi^{0}_{\star})}{f_{X_{0}}(\xi\star)}\\ -\int_{\xi^{0}_{\star}}^{\infty}\frac{v(\xi)}{1-\alpha}\mathrm{d}\xi\end{pmatrix},\overline{\Sigma}\right)\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0,$
where $\overline{\Sigma}$ is given by (14).
Proof 2.5.
@@
$h^{-1}\begin{pmatrix}\xi^{h}_{\star}-\xi^{0}_{\star}\\ \chi^{h}_{\star}-\chi^{0}_{\star}\end{pmatrix}\rightarrow-\begin{pmatrix}\frac{v(\xi^{0}_{\star})}{f_{X_{0}}(\xi\star)}\\ \int_{\xi^{0}_{\star}}^{\infty}\frac{v(\xi)}{1-\alpha}\mathrm{d}\xi\end{pmatrix}\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
Using Theorem 2.2 and the above result concludes the proof.
@@
    Unlike NSA, ANSA scores a convergence speed in $\mathrm{O}(h)$ for both the VaR and ES, for any $\beta\in\big(\frac{1}{2},1\big)$. The choice of $\beta$ in the range $\big(\frac{1}{2},1\big)$ is free of any constraint on the choice of $\gamma_{1}$.
@@
    The covariance matrix (14) is independent of $\gamma_{1}$, reflecting a strong numerical stability for ANSA. The VaR and ES estimates remain nonetheless asymptotically correlated.
@@
    Although the covariance matrix (14) is not computable, it is approximatable. See Remark 1.6(iv) for additional comments as well as an estimation formula of the ES variance factor.
@@
    [9, Theorem 1.3] analyzes a two-time-scale unbiased algorithm where the ES update injects the averaged VaR estimator into (12). We refer to Remarks 1.7(i)-(ii) on the suitability of such result to our biased case. The technique used to prove the aforementioned theorem relies on a spectral analysis of the algorithm’s linearization, coupled with a tightness study of its diagonalized interpolation.
@@ 2.2 Complexity Analysis
achieves a convergence rate in distribution for the ANSA scheme (13)-(8) of order $\varepsilon$ as $\varepsilon\downarrow 0$. The corresponding complexity satisfies
@@
for some positive constant $C<\infty$. By Corollary 2.4, the convergence rate of NSA with bias parameter $h$ and iterations amount $n=\lceil h^{-2}\rceil$ is of order $h$ as $\mathcal{H}\ni h\downarrow 0$, hence the result by setting $h$ such that $h\leq\varepsilon$.
@@
Theorems 1.8 and 2.8 reveal that, unlike NSA, ANSA achieves a complexity of order $\varepsilon^{-3}$ regardless of the value of $\beta\in\big(\frac{1}{2},1\big)$. This allows to circumvent the constraint on $\gamma_{1}$ that appears for the NSA scheme in the optimal case $\beta=1$.
@@ 3 Multilevel Stochastic Approximation Algorithm
$\xi^{h_{L}}_{\star}$ $=\xi^{h_{0}}_{\star}+\sum_{\ell=1}^{L}\xi^{h_{\ell}}_{\star}-\xi^{h_{\ell-1}}_{\star},$ (15)
$\chi^{h_{L}}_{\star}$ $=\chi^{h_{0}}_{\star}+\sum_{\ell=1}^{L}\chi^{h_{\ell}}_{\star}-\chi^{h_{\ell-1}}_{\star}.$ (16)
@@
$\xi^{\text{\tiny\rm ML}}_{\mathbf{N}}$ $=\xi^{h_{0},0}_{N_{0}}+\sum_{\ell=1}^{L}\xi^{h_{\ell},\ell}_{N_{\ell}}-\xi^{h_{\ell-1},\ell}_{N_{\ell}},$ (17)
$\chi^{\text{\tiny\rm ML}}_{\mathbf{N}}$ $=\chi^{h_{0},0}_{N_{0}}+\sum_{\ell=1}^{L}\chi^{h_{\ell},\ell}_{N_{\ell}}-\chi^{h_{\ell-1},\ell}_{N_{\ell}}$. (18)
Each level $0\leq\ell\leq L$ is simulated independently. First, we simulate the level 0 pair $(\xi^{h_{0},0}_{N_{0}},\chi^{h_{0},0}_{N_{0}})$ using $N_{0}$ iterations of the NSA scheme (7)-(8) with bias $h_{0}$. Then, at each level $1\leq\ell\leq L$, given innovations $\{(X^{(n)}_{h_{\ell-1},\ell},X^{(n)}_{h_{\ell},\ell}),1\leq n\leq N_{\ell}\}\mathrel{{\mathop{\sim}\limits^{\text{\rm\tiny i.i.d.}}}}(X_{h_{\ell-1},\ell},X_{h_{\ell},\ell})$, for $j\in\{\ell-1,\ell\}$, we perform $N_{\ell}$ iterations of the dynamics
$\xi^{h_{j},\ell}_{n+1}$ $=\xi^{h_{j},\ell}_{n}-\gamma_{n+1}H_{1}(\xi^{h_{j},\ell}_{n},X_{h_{j},\ell}^{(n+1)}),$ (19)
$\chi^{h_{j},\ell}_{n+1}$ $=\chi^{h_{j},\ell}_{n}-\frac{1}{n+1}H_{2}(\chi^{h_{j},\ell}_{n},\xi^{h_{j},\ell}_{n},X_{h_{j},\ell}^{(n+1)}),$ (20)
@@ 3.1 Convergence Rate Analysis
    $h_{\ell}^{-\frac{1}{2}}(X_{h_{\ell}}-X_{h_{\ell-1}})\eqqcolon G_{\ell}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{\mathcal{L}}}}G\coloneqq\big((M-1)\mathbb{V}\mathrm{ar}(\varphi(Y,Z)|Y)\big)^{\frac{1}{2}}\mathfrak{N}\quad\mbox{ as }\quad\ell\uparrow\infty,$
@@
    Assume that, for all $\ell\geq 1$, $F_{X_{h_{\ell-1}}|G_{\ell}}$ is -$\mathbb{P}\text{\rm-as}$ continuously differentiable with derivative $f_{X_{h_{\ell-1}}|G_{\ell}}$, and that $\{(x,g)\mapstochar\rightarrow f_{X_{h_{\ell-1}}|G_{\ell}=g}(x),\ell\geq 1\}$ are bounded uniformly in $\ell\geq 1$ and converge locally uniformly to some bounded and continuous function $(x,g)\mapstochar\rightarrow f_{g}(x)$. Then, for any $\xi\in\mathbb{R}$,
    ${h_{\ell}^{-\frac{1}{2}}\mathbb{E}\big[\big|\mathds{1}_{X_{h_{\ell}}>\xi}-\mathds{1}_{X_{h_{\ell-1}}>\xi}\big|\big]}\rightarrow\mathbb{E}[|G|f_{G}(\xi)]\quad\mbox{ as }\quad\ell\uparrow\infty.$
@@
    $h_{\ell}^{-\frac{1}{2}}\big((X_{h_{\ell}}-\xi)^{+}-(X_{h_{\ell-1}}-\xi)^{+}\big)\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{\mathcal{L}}}}\mathds{1}_{X_{0}>\xi}\,G\quad\mbox{ as }\quad\ell\uparrow\infty.$
Proof 3.2.
See Appendix D.
Remark $3.$
The framework of Lemma 3.1(ii) strengthens that of [11, Proposition 3.2(ii)] (recalled in Lemma G.1(ii)) by supposing the existence and convergence of pdf for the conditional laws $X_{h_{\ell-1}}\,|\,G_{\ell}$, $\ell\geq 1$. As shown in Step 9 of Theorem 3.5’s proof, Appendix E, such framework allows to retrieve an asymptotic distribution for the martingale part of the VaR error decomposition (77). While remaining weaker than [25, Assumption 2.5], it reaches optimal complexity according to [11, Remark 3.3 & Theorem 3.9(i)].
{assumption}
[[11, Assumption 3.4]] There exist positive constants $C,\delta_{0}<\infty$ such that, for any $h\in\mathcal{H}$ and any compact set $\mathcal{K}\subset\mathbb{R}$,
@@
$N_{\ell}=h_{L}^{-\frac{2}{\beta}}\bigg(\sum_{\ell^{\prime}=0}^{L}h_{\ell^{\prime}}^{-\frac{2\beta-1}{2(1+\beta)}}\bigg)^{\frac{1}{\beta}}h_{\ell}^{\frac{3}{2(1+\beta)}},\quad 0\leq\ell\leq L,$ (21)
@@
$\begin{pmatrix}h_{L}^{-1}\big(\xi^{\text{\tiny\rm ML}}_{\mathbf{N}}-\xi^{h_{L}}_{\star}\big)\\ h_{L}^{-\frac{1}{\beta}-\frac{2\beta-1}{4\beta(1+\beta)}}\big(\chi^{\text{\tiny\rm ML}}_{\mathbf{N}}-\chi^{h_{L}}_{\star}\big)\end{pmatrix}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{\mathcal{L}}}}\mathcal{N}(0,\Sigma^{\text{\tiny\rm ML}}_{\beta})\quad\mbox{ as }\quad L\uparrow\infty,$
@@
$\begin{pmatrix}\frac{\gamma_{1}\mathbb{E}[|G|f_{G}(\xi^{0}_{\star})]}{(1-\alpha)(2f_{X_{0}}(\xi^{0}_{\star})-(1-\alpha)\gamma_{1}^{-1}\mathds{1}_{\beta=1})}&0\\ 0&\frac{h_{0}^{\frac{2\beta-1}{2(1+\beta)}}\big(M^{\frac{2\beta-1}{2(1+\beta)}}-1\big)^{\frac{1}{\beta}}}{(1-\alpha)^{2}}\Big(\frac{h_{0}^{-1}\mathbb{V}\mathrm{ar}((X_{h_{0}}-\xi^{h_{0}}_{\star})^{+})}{M^{\frac{2\beta-1}{2\beta(1+\beta)}}}+\frac{\mathbb{V}\mathrm{ar}(\mathds{1}_{X_{0}>\xi^{0}_{\star}}\,G)}{M^{\frac{2\beta-1}{2(1+\beta)}}-1}\Big)\end{pmatrix}$, (22)
@@
    We used in (21) the optimal iterations amounts for the VaR MLSA scheme according to [11, Theorem 3.9(i)]. They were obtained by optimizing the MLSA complexity while constraining the VaR estimation error to a prescribed accuracy $\varepsilon$.
@@
    The VaR and ES evolve at different speeds: in $\mathrm{O}(h_{L})$ for the VaR and in $\mathrm{O}\big(h_{L}^{\frac{1}{\beta}+\frac{2\beta-1}{4\beta(1+\beta)}}\big)$ for the ES as $L\uparrow\infty$. The optimal ES convergence speed in $\mathrm{O}\big(h_{L}^{\frac{9}{8}}\big)$ is attained for $\beta=1$ under the condition $\lambda\gamma_{1}>1$.
@@
    The VaR variance factor depends on $\gamma_{1}$ in the same way the NSA covariance matrix (10) does (Remark 1.1(ii)).
    To ensure the additional condition $\lambda\gamma_{1}>1$ in the optimal case $\beta=1$, a rule of thumb would be to choose $\gamma_{1}$ large enough with regard to the variables of the problem. But this is not satisfactory, inasmuch as the VaR component of the covariance matrix $\Sigma_{1}$ in (10) evolves in $\mathrm{O}(\gamma_{1})$ for $\gamma_{1}$ large, damaging the algorithm’s convergence.
    Setting it too small blows up the VaR convergence factor, as it evolves in $\mathrm{O}\big(\big(\gamma_{1}-\frac{1-\alpha}{2f_{X_{0}}(\xi^{0}_{\star})}\big)^{-1}\big)$ if $\gamma_{1}\downarrow\frac{1-\alpha}{2f_{X_{0}}(\xi^{0}_{\star})}$.
@@
    Setting $h_{0}$ sufficiently small should help reduce $\mathbb{V}\mathrm{ar}((X_{h_{0}}-\xi^{h_{0}}_{\star})^{+})$, hence the impact of the initial level simulation on the ES variance. However, as clarified in Theorem 3.8, for a prescribed accuracy $\varepsilon\in(0,1)$, $h_{0}$ must be floored by $\varepsilon$, and $L$ set as in (25). Reducing $h_{0}$ would consequently reduce $L$, hence exiting the asymptotic regime. $h_{0}$ should therefore be fine-tuned. See [18, Remark 2.6] for further comments on setting $h_{0}$.
@@
    The VaR-ES asymptotic MLSA correlation is null, which is linked to the speed differential between the VaR and ES. It expresses an independent behavior of the VaR and ES as $L\uparrow\infty$. The strong stability of the ES is thus hardly affected by the instability of the VaR. This may seem surprising since the scheme (20) clearly incorporates the VaR estimates in the ES approximation. However, it can be explained by the robustifying effect of the averaging property of the ES estimates (20).

### v5

@@ 1.2 Complexity Analysis
achieves a convergence rate in distribution for the NSA scheme (1.4)-(1.5) of order $\varepsilon$ as $\varepsilon\downarrow 0$. The corresponding complexity satisfies
@@
for some positive constant $C<\infty$. By Corollary 1.5, the convergence rate of NSA, with bias parameter $h$ and iterations amount $n=\lceil h^{-2}\rceil$, is of order $h^{\beta}$ as $\mathcal{H}\ni h\downarrow 0$. The result follows by setting $h$ such that $h^{\beta}\leq\varepsilon$. ∎
@@ 2 Averaged Nested Stochastic Approximation Algorithm
According to Theorem 1.4 and Corollary 1.5, the best CLT convergence rate is achieved for $\beta=1$, i.e. by setting $\gamma_{n}=\gamma_{1}n^{-1}$ for the VaR NSA scheme. This choice is only possible under the additional constraint $\lambda\gamma_{1}>1$, where $\lambda$ is an explicit yet inaccessible constant (1.6). In practice, $\gamma_{1}$ must be fine-tuned, adding a computational burden to the algorithm’s usage.
@@
$\overline{\xi}^{h}_{n}=\frac{1}{n}\sum_{k=1}^{n}\xi^{h}_{k}=\Big(1-\frac{1}{n}\Big)\overline{\xi}^{h}_{n-1}+\frac{1}{n}\xi^{h}_{n},\quad n\geq 1,$ (2.1)
where $\overline{\xi}^{h}_{0}=0$ and $\{\xi^{h}_{n},n\geq 0\}$ are obtained via the NSA scheme (1.4) with step sizes $\{\gamma_{n}=\gamma_{1}n^{-\beta},n\geq 1\}$, $\gamma_{1}>0$, $\beta\in\big(\frac{1}{2},1\big)$.
@@ 2.1 Convergence Rate Analysis
$h^{-1}\begin{pmatrix}\overline{\xi}^{h}_{\lceil h^{-2}\rceil}-\xi^{h}_{\star}\\ \chi^{h}_{\lceil h^{-2}\rceil}-\chi^{h}_{\star}\end{pmatrix}\mathrel{{\mathop{\longrightarrow}\limits^{\mathcal{L}}}}\mathcal{N}(0,\overline{\Sigma})\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0,$
@@
$\overline{\Sigma}=\begin{pmatrix}\frac{\alpha(1-\alpha)}{f_{X_{0}}(\xi^{0}_{\star})^{2}}&\frac{\alpha}{1-\alpha}\frac{\mathbb{E}[(X_{0}-\xi^{0}_{\star})^{+}]}{f_{X_{0}}(\xi^{0}_{\star})}\\ \frac{\alpha}{1-\alpha}\frac{\mathbb{E}[(X_{0}-\xi^{0}_{\star})^{+}]}{f_{X_{0}}(\xi^{0}_{\star})}&\frac{\operatorname{\mathbb{V}ar}((X_{0}-\xi^{0}_{\star})^{+})}{(1-\alpha)^{2}}\end{pmatrix}$. (2.2)
Proof.
See Appendix C. ∎
@@
$h^{-1}\begin{pmatrix}\overline{\xi}^{h}_{\lceil h^{-2}\rceil}-\xi^{0}_{\star}\\ \chi^{h}_{\lceil h^{-2}\rceil}-\chi^{0}_{\star}\end{pmatrix}\mathrel{{\mathop{\longrightarrow}\limits^{\mathcal{L}}}}\mathcal{N}\left(\begin{pmatrix}-\frac{v(\xi^{0}_{\star})}{f_{X_{0}}(\xi\star)}\\ -\int_{\xi^{0}_{\star}}^{\infty}\frac{v(\xi)}{1-\alpha}\mathrm{d}\xi\end{pmatrix},\overline{\Sigma}\right)\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0,$
where $\overline{\Sigma}$ is given by (2.2).
Proof.
@@
$h^{-1}\begin{pmatrix}\xi^{h}_{\star}-\xi^{0}_{\star}\\ \chi^{h}_{\star}-\chi^{0}_{\star}\end{pmatrix}\to-\begin{pmatrix}\frac{v(\xi^{0}_{\star})}{f_{X_{0}}(\xi\star)}\\ \int_{\xi^{0}_{\star}}^{\infty}\frac{v(\xi)}{1-\alpha}\mathrm{d}\xi\end{pmatrix}\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
Using Theorem 2.2 and the above result concludes the proof. ∎
@@
    Unlike NSA, ANSA scores a convergence speed in $\operatorname{O}(h)$ for both the VaR and ES, for any $\beta\in\big(\frac{1}{2},1\big)$. The choice of $\beta$ in the range $\big(\frac{1}{2},1\big)$ is free of any constraint on the choice of $\gamma_{1}$.
@@
    The covariance matrix (2.2) is independent of $\gamma_{1}$, reflecting a strong numerical stability for ANSA. The VaR and ES estimates remain nonetheless asymptotically correlated.
@@
    Although the covariance matrix (2.2) is not computable, it is approximatable. See Remark 1.6(iv) for additional comments as well as an estimation formula of the ES variance factor.
@@
    [9, Theorem 1.3] analyzes a two-time-scale unbiased algorithm where the ES update injects the averaged VaR estimator into (1.9). We refer to Remarks 1.7(i)-(ii) on the suitability of such result to our biased case. The technique used to prove the aforementioned theorem relies on a spectral analysis of the algorithm’s linearization, coupled with a tightness study of its diagonalized interpolation.
@@ 2.2 Complexity Analysis
achieves a convergence rate in distribution for the ANSA scheme (2.1)-(1.5) of order $\varepsilon$ as $\varepsilon\downarrow 0$. The corresponding complexity satisfies
@@
for some positive constant $C<\infty$. By Corollary 2.3, the convergence rate of NSA with bias parameter $h$ and iterations amount $n=\lceil h^{-2}\rceil$ is of order $h$ as $\mathcal{H}\ni h\downarrow 0$, hence the result by setting $h$ such that $h\leq\varepsilon$. ∎
@@
Theorems 1.8 and 2.6 reveal that, unlike NSA, ANSA achieves a complexity of order $\varepsilon^{-3}$ regardless of the value of $\beta\in\big(\frac{1}{2},1\big)$. This allows to circumvent the constraint on $\gamma_{1}$ that appears for the NSA scheme in the optimal case $\beta=1$.
@@ 3 Multilevel Stochastic Approximation Algorithm
$\xi^{h_{L}}_{\star}$ $=\xi^{h_{0}}_{\star}+\sum_{\ell=1}^{L}\xi^{h_{\ell}}_{\star}-\xi^{h_{\ell-1}}_{\star},$ (3.1)
$\chi^{h_{L}}_{\star}$ $=\chi^{h_{0}}_{\star}+\sum_{\ell=1}^{L}\chi^{h_{\ell}}_{\star}-\chi^{h_{\ell-1}}_{\star}.$ (3.2)
@@
$\xi^{\text{\tiny\rm ML}}_{\mathbf{N}}$ $=\xi^{h_{0},0}_{N_{0}}+\sum_{\ell=1}^{L}\xi^{h_{\ell},\ell}_{N_{\ell}}-\xi^{h_{\ell-1},\ell}_{N_{\ell}},$ (3.3)
$\chi^{\text{\tiny\rm ML}}_{\mathbf{N}}$ $=\chi^{h_{0},0}_{N_{0}}+\sum_{\ell=1}^{L}\chi^{h_{\ell},\ell}_{N_{\ell}}-\chi^{h_{\ell-1},\ell}_{N_{\ell}}$. (3.4)
Each level $0\leq\ell\leq L$ is simulated independently. First, we simulate the level 0 pair $(\xi^{h_{0},0}_{N_{0}},\chi^{h_{0},0}_{N_{0}})$ using $N_{0}$ iterations of the NSA scheme (1.4)-(1.5) with bias $h_{0}$. Then, at each level $1\leq\ell\leq L$, given innovations $\{(X^{(n)}_{h_{\ell-1},\ell},X^{(n)}_{h_{\ell},\ell}),1\leq n\leq N_{\ell}\}\mathrel{{\mathop{\sim}\limits^{\text{\rm\tiny i.i.d.}}}}(X_{h_{\ell-1},\ell},X_{h_{\ell},\ell})$, for $j\in\{\ell-1,\ell\}$, we perform $N_{\ell}$ iterations of the dynamics
$\xi^{h_{j},\ell}_{n+1}$ $=\xi^{h_{j},\ell}_{n}-\gamma_{n+1}H_{1}(\xi^{h_{j},\ell}_{n},X_{h_{j},\ell}^{(n+1)}),$ (3.5)
$\chi^{h_{j},\ell}_{n+1}$ $=\chi^{h_{j},\ell}_{n}-\frac{1}{n+1}H_{2}(\chi^{h_{j},\ell}_{n},\xi^{h_{j},\ell}_{n},X_{h_{j},\ell}^{(n+1)}),$ (3.6)
@@ 3.1 Convergence Rate Analysis
    $h_{\ell}^{-\frac{1}{2}}(X_{h_{\ell}}-X_{h_{\ell-1}})\eqqcolon G_{\ell}\mathrel{{\mathop{\longrightarrow}\limits^{\mathcal{L}}}}G\coloneqq\big((M-1)\operatorname{\mathbb{V}ar}(\varphi(Y,Z)|Y)\big)^{\frac{1}{2}}\mathfrak{N}\quad\mbox{ as }\quad\ell\uparrow\infty,$
@@
    Assume that, for all $\ell\geq 1$, $F_{X_{h_{\ell-1}}|G_{\ell}}$ is -$\operatorname{\mathds{P}\text{-as}}$ continuously differentiable with derivative $f_{X_{h_{\ell-1}}|G_{\ell}}$, and that $\{(x,g)\mapsto f_{X_{h_{\ell-1}}|G_{\ell}=g}(x),\ell\geq 1\}$ are bounded uniformly in $\ell\geq 1$ and converge locally uniformly to some bounded and continuous function $(x,g)\mapsto f_{g}(x)$. Then, for any $\xi\in\mathbb{R}$,
    ${h_{\ell}^{-\frac{1}{2}}\mathbb{E}\big[\big|\mathds{1}_{X_{h_{\ell}}>\xi}-\mathds{1}_{X_{h_{\ell-1}}>\xi}\big|\big]}\to\mathbb{E}[|G|f_{G}(\xi)]\quad\mbox{ as }\quad\ell\uparrow\infty.$
@@
    $h_{\ell}^{-\frac{1}{2}}\big((X_{h_{\ell}}-\xi)^{+}-(X_{h_{\ell-1}}-\xi)^{+}\big)\mathrel{{\mathop{\longrightarrow}\limits^{\mathcal{L}}}}\mathds{1}_{X_{0}>\xi}\,G\quad\mbox{ as }\quad\ell\uparrow\infty.$
Proof.
See Appendix D. ∎
Remark 3.2.
The framework of Lemma 3.1(ii) strengthens that of [11, Proposition 3.2(ii)] (recalled in Lemma G.1(ii)) by supposing the existence and convergence of pdf for the conditional laws $X_{h_{\ell-1}}\,|\,G_{\ell}$, $\ell\geq 1$. As shown in Step 9 of Theorem 3.5’s proof, Appendix E, such framework allows to retrieve an asymptotic distribution for the martingale part of the VaR error decomposition (E.1). While remaining weaker than [25, Assumption 2.5], it reaches optimal complexity according to [11, Remark 3.3 & Theorem 3.9(i)].
Assumption 3.3 ([11, Assumption 3.4]).
There exist positive constants $C,\delta_{0}<\infty$ such that, for any $h\in\mathcal{H}$ and any compact set $\mathcal{K}\subset\mathbb{R}$,
@@
$N_{\ell}=h_{L}^{-\frac{2}{\beta}}\bigg(\sum_{\ell^{\prime}=0}^{L}h_{\ell^{\prime}}^{-\frac{2\beta-1}{2(1+\beta)}}\bigg)^{\frac{1}{\beta}}h_{\ell}^{\frac{3}{2(1+\beta)}},\quad 0\leq\ell\leq L,$ (3.7)
@@
$\begin{pmatrix}h_{L}^{-1}\big(\xi^{\text{\tiny\rm ML}}_{\mathbf{N}}-\xi^{h_{L}}_{\star}\big)\\ h_{L}^{-\frac{1}{\beta}-\frac{2\beta-1}{4\beta(1+\beta)}}\big(\chi^{\text{\tiny\rm ML}}_{\mathbf{N}}-\chi^{h_{L}}_{\star}\big)\end{pmatrix}\mathrel{{\mathop{\longrightarrow}\limits^{\mathcal{L}}}}\mathcal{N}(0,\Sigma^{\text{\tiny\rm ML}}_{\beta})\quad\mbox{ as }\quad L\uparrow\infty,$
@@
$\begin{pmatrix}\frac{\gamma_{1}\mathbb{E}[|G|f_{G}(\xi^{0}_{\star})]}{(1-\alpha)(2f_{X_{0}}(\xi^{0}_{\star})-(1-\alpha)\gamma_{1}^{-1}\mathds{1}_{\beta=1})}&0\\ 0&\frac{h_{0}^{\frac{2\beta-1}{2(1+\beta)}}\big(M^{\frac{2\beta-1}{2(1+\beta)}}-1\big)^{\frac{1}{\beta}}}{(1-\alpha)^{2}}\Big(\frac{h_{0}^{-1}\operatorname{\mathbb{V}ar}((X_{h_{0}}-\xi^{h_{0}}_{\star})^{+})}{M^{\frac{2\beta-1}{2\beta(1+\beta)}}}+\frac{\operatorname{\mathbb{V}ar}(\mathds{1}_{X_{0}>\xi^{0}_{\star}}\,G)}{M^{\frac{2\beta-1}{2(1+\beta)}}-1}\Big)\end{pmatrix}$, (3.8)
@@
    We used in (3.7) the optimal iterations amounts for the VaR MLSA scheme according to [11, Theorem 3.9(i)]. They were obtained by optimizing the MLSA complexity while constraining the VaR estimation error to a prescribed accuracy $\varepsilon$.
@@
    The VaR and ES evolve at different speeds: in $\operatorname{O}(h_{L})$ for the VaR and in $\operatorname{O}\big(h_{L}^{\frac{1}{\beta}+\frac{2\beta-1}{4\beta(1+\beta)}}\big)$ for the ES as $L\uparrow\infty$. The optimal ES convergence speed in $\operatorname{O}\big(h_{L}^{\frac{9}{8}}\big)$ is attained for $\beta=1$ under the condition $\lambda\gamma_{1}>1$.
@@
    The VaR variance factor depends on $\gamma_{1}$ in the same way the NSA covariance matrix (1.7) does (Remark 1.2(ii)).
    To ensure the additional condition $\lambda\gamma_{1}>1$ in the optimal case $\beta=1$, a rule of thumb would be to choose $\gamma_{1}$ large enough with regard to the variables of the problem. But this is not satisfactory, inasmuch as the VaR component of the covariance matrix $\Sigma_{1}$ in (1.7) evolves in $\operatorname{O}(\gamma_{1})$ for $\gamma_{1}$ large, damaging the algorithm’s convergence.
    Setting it too small blows up the VaR convergence factor, as it evolves in $\operatorname{O}\big(\big(\gamma_{1}-\frac{1-\alpha}{2f_{X_{0}}(\xi^{0}_{\star})}\big)^{-1}\big)$ if $\gamma_{1}\downarrow\frac{1-\alpha}{2f_{X_{0}}(\xi^{0}_{\star})}$.
@@
    Setting $h_{0}$ sufficiently small should help reduce $\operatorname{\mathbb{V}ar}((X_{h_{0}}-\xi^{h_{0}}_{\star})^{+})$, hence the impact of the initial level simulation on the ES variance. However, as clarified in Theorem 3.7, for a prescribed accuracy $\varepsilon\in(0,1)$, $h_{0}$ must be floored by $\varepsilon$, and $L$ set as in (3.11). Reducing $h_{0}$ would consequently reduce $L$, hence exiting the asymptotic regime. $h_{0}$ should therefore be fine-tuned. See [18, Remark 2.6] for further comments on setting $h_{0}$.
@@
    The VaR-ES asymptotic MLSA correlation is null, which is linked to the speed differential between the VaR and ES. It expresses an independent behavior of the VaR and ES as $L\uparrow\infty$. The strong stability of the ES is thus hardly affected by the instability of the VaR. This may seem surprising since the scheme (3.6) clearly incorporates the VaR estimates in the ES approximation. However, it can be explained by the robustifying effect of the averaging property of the ES estimates (3.6).